OpenStudy (anonymous):

The slope of the line perpendicular to y = 3/4 x - 7 is A. 3/4 B. -4/3 C. 4/3 D. -3/4

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

Y=aX+b a is the slope, so the answer is A

OpenStudy (anonymous):

b

OpenStudy (anonymous):

@andy: she doesnt want the slope of this line but the slope of a line perpendicular to it (ap) which is with your equation y=a*x+b ap=-1/a

OpenStudy (anonymous):

b. the reason it is not a. is because parallel lines have the same gradient, hence a parallel line would have slope 3/4 (the number in front of the x). Perpendicular lines have -1/m gradient, in your case m=3/4. So the perpendicular line has -1/(3/4) =-4/3. Answer b.

OpenStudy (anonymous):

hmmm